OpenStudy (anonymous):

21 is 35% of what number? A. 21=0.35x;62.4 B. x=21x0.35;7.4 C. 0.35=21X;40x8 D. 21=0.35x;60

OpenStudy (anonymous):

21*100/35 =60

OpenStudy (anonymous):

Let that "some number" be x. Then we can rewrite that sentence as, 21 = 0.35x Dividing both sides by 0.35 we arrive at the solution: x = 60

OpenStudy (anonymous):

Or use a proportion\[35/100=21/x \rightarrow 35x=2100\rightarrow x=60\]